OpenStudy (anonymous):

If p and r are integers, p does not equal 0, and p = -r, which of the following must be true? a. p[]p>r b. p+r<0 c. p-r<0 d. pr<0

OpenStudy (blurbendy):

what does a say? im not familiar with what [] means

OpenStudy (anonymous):

I honestly don't know its just an empty box

OpenStudy (blurbendy):

really?? that's so weird

OpenStudy (anonymous):

yea I know and I don't know how to answer it

OpenStudy (blurbendy):

well if you let p = -7 and r = 7 this fits the definition, because p = -r so let's do the last 3 -7 + 7 = 0 < 0 is false -7 - 7 = -14 < 0 is true 7*-7 = -49 < 0 is true

OpenStudy (blurbendy):

oops, -7 - (-7) = -7 + 7 < 0 is false

OpenStudy (blurbendy):

only the last one is true

OpenStudy (anonymous):

thank you

OpenStudy (blurbendy):

yw